HUBBLE SPACE TELESCOPE DAILY REPORT #3072 PERIOD COVERED: 0000Z (UTC) 03/13/02 - 0000Z (UTC) 03/14/02 Daily Status Report as of 073/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed HST 8968 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} was used to characterize uncompensated environmental disturbances acting upon the HST during normal operation. The VDT is a passive test {not a forced-response test} used to obtain signatures for both externally induced {e.g. SA-3} and internally induced {e.g. NCC and ACS mechanisms} disturbances for comparison with past VDT results. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. The test occurs after release, and most of the VDT can be run during the BEA period. The -V1 sunpoint portion of the VDT occurs after the BEA period is complete. There were no reported problems. 1.2 Completed ACS/WFC 9006 (CCD Temperature Set Point Determination) The Advanced Camera for Surveys (WFC) was used to adjust the CCD temperature to nominal operating level. The proposal completed nominally. 1.3 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.4 Completed STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.5 Completed STIS/MA1/MA2 8959 (MAMA Dark vs. Temperature) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to measure the dark rate for both MAMA detectors as a function of instantaneous temperature and recent thermal history. These will be used to make measurements at the beginning and end of three consecutive SAA-free orbit blocks. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 2 Successful: 2 Scheduled Re-acquisitions: 1 Successful: 1 2.2 FHST Updates: Scheduled: 6 Successful: 6 2.3 Operations Notes: As directed by operations notes, operations personnel made limit changes for the NICMOS Cooling system and modified EPS limits to post-SM3B values. Throughout the day, three gyro bias updates and three attitude reference updates were uplinked (ROPs DF-14A and PO-5). Per HSTAR 8550, there were bit flips in three of the dumps performed in accordance with ACS proposal 9003, performed yesterday. The Vehicle Disturbance Test(VDT) #1 (see description in 2.1) executed nominally last night between 072/003000Z and 072/070000Z. VDT #1 exercised ACS mechanisms: Cal Door, WFC and HRC shutters, the SBC filter wheel, and the two WFC/HRC filter wheels. ACS mechanisms were articulated every 4-minutes, and each mechanism was cycled three times during the 2-1/2 hour test. All mechanism motions executed nominally. VDT #1 was the first look at the HST disturbance environment using maneuver gains since the post-SM3A VDT, and the good news is that HST disturbances are now 20-times lower than disturbances observed with SA1/SA2 present. Previously, terminator disturbance amplitudes reached 1.0 arcsec prior to SM3B, whereas now the maximum disturbances observed are due to reaction wheel zero-speed crossings, approximately 50 mas in amplitude. HST jitter was observed during ACS mechanism articulation, and the largest mechanism disturbances were ~25 mas during motion of the WFC/HRC filter wheels. HST jitter using nominal control law gains will be approximately 10-times lower than the peak jitter observed during the VDT, so peak ACS filter wheel disturbances are expected to be ~2.5 mas, buried in the rate gyro noise. VDT#2 will be conducted this afternoon. This test will establish a jitter baseline signature prior to activating the NICMOS Cooling System. A TTR was written for a generic late TDRSS acquisition at 073/0634Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW